OpenStudy (anonymous):

Will someone help? Choose the equivalent system of linear equations that will produce the same solution as the one given below. 6x + 2y = -6 3x - 4y = -18 12x + 4y = -12 15x = -30 8x + 4y = -4 14x = -10 6x - 8y = -36 -6y = -42 6x - y = -15 3y = 9

OpenStudy (anonymous):

@DanJS

OpenStudy (anonymous):

I have this question and then two others just like it.... I have been working on this assignment for a very long time and I would just like to get it done. Do you think you could help me finish? please..

OpenStudy (danjs):

For this one, Try multiplying the first equation by 2

OpenStudy (danjs):

you can multiply an equation through by a number on both sides, and it will not change the equation at all overall

OpenStudy (danjs):

6x + 2y = -6 Multiply everything by 2 12x + 4y = -12

OpenStudy (anonymous):

6x + 2y = -6 12x +4y= Like that...

OpenStudy (danjs):

yea, but you have to do it to both sides, so that the equation doesnt change

OpenStudy (anonymous):

Well I forgot the -6 but

OpenStudy (anonymous):

So 12x+4y=-12

OpenStudy (danjs):

12x + 4y = -12 3x - 4y = -18 Those are the two equations now right? i just put the second one back below the first

OpenStudy (danjs):

Now you need to add the two equations together

OpenStudy (danjs):

12x + 4y = -12 3x - 4y = -18 ---------------- 15x + 0y = -30

OpenStudy (danjs):

so you get 15x = -30

OpenStudy (danjs):

That is what they are looking for, 12x + 4y = -12 and 15x = -30

OpenStudy (anonymous):

I have to go but I might be able to long on either later today or tomorrow. Can we just finish it later please? @DanJS

OpenStudy (anonymous):

Log*

OpenStudy (danjs):

sure